# PipelineRL: Faster On-policy Reinforcement Learning for Long Sequence Generation

### v1

@@ 2.1 Reinforcement Learning for Large Language Models
In most practical RL setups, the current policy $\pi$ will often differ from the behavior policy $\mu$ that generates $y_{k}$, due to the weights lagging, quantization or implementation difference between the inference and training softwares. This difference is usually handled by either a trust region constraint (Schulman et al., 2017) or using Importance Sampling (IS). In practice, the importance weights are truncated to reduce the variance of the estimator (Munos et al., 2016; Espeholt et al., 2018):
@@ 5 Experiments
For each experiment, we train the Qwen 2.5 base model (Yang et al., 2024) with 7B parameters on 17K math problems from the OpenReasoner Zero dataset (Hu et al., 2025) for 1000 optimizer steps with the batch size $B=1024$. We use Adam optimizer (Kingma, 2014) with the learning rate 1e-6. We run the PipelineRL experiments on 16 DGX-H100 nodes, using 48 GPUs for generation at batch size $H=64$ and 80 GPUs for training. We tweak PipelineRL to simulate Conventional RL by accumulating and shuffling a buffer of $BG$ samples at the Preprocessor stage before the $G$ optimizer steps of each RL step start. To estimate the Conventional RL throughput, we use 2 nodes for generation at batch size $H=64$ and 2 nodes for training, and then add a correction for training on 8x fewer GPUs than what an efficient Conventional RL implementation with a quick generation-training transition could use. To estimate the inference throughput on 128 GPUS instead of 16 GPUs, we submit $128/16$ batches of $16\times G\times 1024/128$ and take the maximum completion time. We give reward 1 to any generated sequence with the correct answer and 0 otherwise. We also give a soft penalty to the model when it gets close to the max sequence length. We train every model with importance weighted REINFORCE as described in Section 2 and clamp the importance weights to 5. For our experiment we use vLLM (Kwon et al., 2023a) to generate trajectories and use DeepSpeed (Rasley et al., 2020) through accelerate to train our model.
@@ 5.1 Impact of in-flight weight updates on on-policyness
In this section, our goal is to investigate whether mixed-policy sequences, produced by in-flight weight updates, curtail on-policyness in Pipeline-RL. To this end, we generate sequences under a range of conditions based on the recency of the policy used during token generation: (1) mixed-policy tokens with KV cache reuse, as in our Pipeline-RL implementation, (2) mixed-policy tokens with no KV cache reuse, i.e. recomputing attention scores for all tokens, (3) fully off-policy tokens. We measure KL-divergence between the distribution of fully on-policy sequences with the other sequences. We quantify the in-flight impacts by measuring the KL divergence between the distribution of fully on-policy sequences and other sequence types. Ideally, the mixed-policy distributions should closely align the fully on-policy distribution in order to have .
We design a controlled setup to generate different types of sequences. For mixed-policy tokens, we first run Pipeline-RL training and save model checkpoints at every optimizer step. Then, starting from a checkpoint at timestep $C$, we sample sequences of length $\frac{L}{g_{\max}}$ tokens where L is the maximum sequence length and $g_{\max}$ is the maximum lag. This approach simulates in-flight weight updates as in Pipeline-RL. Here, we make a simplifying assumption that in-flight weights occur uniformly across tokens. At each iteration $i$, we sample sequences of $\frac{L}{g_{\max}}$ tokens with the model checkpoint from timestep $C+i$. The generation process continues until either the maximum lag $g_{\max}$ is reached or an EOS token is generated. Two options are considered for mixed-policy tokens, whether we retain KV cache from previous generation steps or we discard them and recompute attention scores. Fully off-policy tokens are when a sequence is generated fully by the checkpoint at timestep $C$, while the current optimizer step is $C+g_{\max}$. Likewise, fully on-policy tokens correspond to generating a sequence completely by the checkpoint $C+g_{\max}$.
In this experiment, we fine-tune Qwen 2.5 7B (base) on the OpenReasoner Zero (Hu et al., 2025) data for 232 optimizer steps. We consider three stages in training to measure KL-divergence: starting at checkpoint 0, 100, and 200. The maximum lag $G$ is set to 32. As presented in Figure˜7, the distribution of mixed-policy sequences closely aligns with that of fully on-policy sequences across all three stages in the training. In contrast, off-policy sequences exhibit consistently higher divergences as lag increases. Also, using stale KV-cache for mixed policy sequences introduces only slightly higher divergence compared to recomputing the cache. This supports our design choice in Pipeline-RL to opt for the more efficient approach of retaining the KV cache.
Figure 7: Impact of mixed-policy sequences on on-policyness. The distribution of mixed-policy tokens is close to that of fully on-policy tokens while the distribution of off-policy tokens exhibit high divergence. These results are consistent across three stages of training where we take a checkpoint at optimizer step $C$ as a reference and compute KL-divergence of subsequent steps with respect to it. Also, reusing stale KV cache does not result in high divergence, thus corroborating our design deicision in Pipeline-RL.
@@ 6 Related work
There exist several other scalable open-source RL implementations. veRL (Sheng et al., 2024) implements Conventional RL efficiently by using a sophisticated hybrid generation-training engine that supports quick transitions between training and generation on the same GPUs. We believe veRL’s throughput would be similar to our Conventional RL baseline. Without the hybrid engine, in OpenRLHF  (Hu et al., 2024) training GPUs idle during generation and vice-versa.
@@ A.2 Conventional RL throughput
where $h(l)$ is the number of sequences still in progress after $l$ steps of decoding, and $U(h)$ is the GPU utilization at batch size $h$. To understand Equation˜10, recall that generating $k$ tokens by definitions takes $k$ flashes under perfect GPU utilization and $k/U(k)$ at the utilization $U(k)$.

### v2

@@ 2.1 Reinforcement Learning for Large Language Models
In most practical RL setups, the current policy $\pi$ will often differ from the behavior policy $\mu$ that generates $y_{k}$, due to the weights lagging, quantization or implementation difference between the inference and training softwares. This difference is usually handled by either a trust region constraint (Schulman et al., 2017) or using Importance Sampling (IS). In practice, the importance sampling weights are truncated to reduce the variance of the estimator (Munos et al., 2016; Espeholt et al., 2018):
@@ 5 Experiments
For each experiment, we train the Qwen 2.5 base model (Yang et al., 2024) with 7B parameters on 17K math problems from the OpenReasoner Zero dataset (Hu et al., 2025) for 1000 optimizer steps with the batch size $B=1024$. We use Adam optimizer (Kingma, 2014) with the learning rate 1e-6. We run the PipelineRL experiments on 16 DGX-H100 nodes, using 48 GPUs for generation at batch size $H=64$ and 80 GPUs for training. We tweak PipelineRL to simulate Conventional RL by accumulating and shuffling a buffer of $BG$ samples at the Preprocessor stage before the $G$ optimizer steps of each RL step start. To estimate the Conventional RL throughput, we use 2 nodes for generation at batch size $H=64$ and 2 nodes for training, and then add a correction for training on 8x fewer GPUs than what an efficient Conventional RL implementation with a quick generation-training transition could use. To estimate the inference throughput on 128 GPUS instead of 16 GPUs, we submit $128/16$ batches of $\tfrac{16\times 1024\times G}{128}$ and take the maximum completion time. We give reward 1 to any generated sequence with the correct answer and 0 otherwise. We also give a soft penalty to the model when it gets close to the max sequence length. We train every model with importance weighted REINFORCE as described in Section 2 and clamp the importance weights to 5. For our experiment we use vLLM (Kwon et al., 2023a) to generate trajectories and use DeepSpeed (Rasley et al., 2020) through accelerate to train the model.
@@ 5.1 Impact of in-flight weight updates on on-policyness
In this section, we compare the sampling distribution of in-flight weight updates to 1) conventional RL with different max lag and 2) in-flight weight update with KV cache recomputation. For this experiment, we save a set of consecutive checkpoints $C_{i}$, one after every optimizer step. To replicate the in-flight weight update, we start from a checkpoint and update the weights of the behavior policy every $\frac{L}{g_{\max}}$ tokens with the subsequent checkpoint, where L is the maximum sequence length and $g_{\max}$ is the maximum lag. Specifically, the PipelineRL behavior policy is defined as:
$\mu:=\mu_{C}(x_{1:t_{1}})\ldots\mu_{C+g}(x_{t_{g}:t_{g+1}}\mid\hat{x}_{1:t_{1}},\ldots\hat{x}_{t_{g-1}:t_{g}})$ (8)
where $t_{1}=\tfrac{2L}{g_{\max}}$ and $t_{g}=t_{g-1}+\tfrac{L}{g_{\max}}$ tokens for lag $g>1$ since the first weight update takes longer than the next updates due to the bubble at the beginning of training, see Figure˜1 b). We also use $\hat{x}$ to stress that the KV cache for the previous tokens is stale - as it was computed under previous model weights. We then compute the KL between the mixed behavior policy $\mu_{C:C+g}$ and the on-policy behavior policy $\mu_{C+g}$. We also report the KL with the mixed behavior policy with updated KV cache which we denote as PipelineRL with KV cache recomputed. To replicate conventional RL, we sample $N$ sequences from the behavior policy $\mu:=\mu_{C}$ and compute the KL with on-policy behavior policy $\mu_{C+g}$ for different lag $k$.
In this experiment, we fine-tune Qwen 2.5 base 7B on the OpenReasoner Zero (Hu et al., 2025) data for 222 optimizer steps. We consider three stages in training to measure KL-divergence: starting at checkpoint 0, 100, and 190. The maximum lag $g_{\max}$ is set to 32 and the maximum sequence length $L$ is 2048. As presented in Figure˜7, the distribution of mixed-policy sequences closely aligns with that of fully on-policy sequences across all three stages in the training. In contrast, off-policy sequences exhibit consistently higher divergences as lag increases. Also, using stale KV-cache for mixed policy sequences introduces only slightly higher divergence compared to recomputing the cache. This supports our design choice in Pipeline-RL to opt for the more efficient approach of retaining the KV cache.
Figure 7: For three different starting checkpoint, PipelineRL with and without KV cache recomputation stay more on-policy than Conventional RL as measured by the KL divergence.
@@ 6 Related work
There exist several other scalable open-source RL implementations. veRL (Sheng et al., 2024) implements Conventional RL efficiently by using a sophisticated hybrid generation-training engine that supports quick transitions between training and generation on the same GPUs. We believe veRL’s throughput would be similar to our Conventional RL baseline. Without the hybrid engine, in OpenRLHF  (Hu et al., 2024) training GPUs idle during generation and vice-versa. Concurrently, Magistral(Mistral-AI et al., 2025) also introduced in-flight weight updates.
@@ A.2 Conventional RL throughput
where $h(l)$ is the number of sequences still in progress after $l$ steps of decoding, and $U(h)$ is the GPU utilization at batch size $h$. To understand Equation˜11, recall that generating $k$ tokens by definitions takes $k$ flashes under perfect GPU utilization and $k/U(k)$ at the utilization $U(k)$.